# Narratives of Collective Action in YouTube’s Discourse on Veganism

### v1

@@ Introduction
Storytelling is a cornerstone of modern communication, with its influence permeating diverse realms from marketing to political discourse (Lund, Cohen, and Scarles 2018; Seargeant 2020). The power of compelling narratives extends beyond individual persuasion and holds the potential to catalyze collective shifts in opinion, mobilize consensus, and foster social cooperation toward shared goals. Social media have democratised opportunities for creating and sharing narratives on a global scale, empowering individuals to shape opinion formation, democratic deliberation, and collective action (Yasseri et al. 2016; Jennings, Suzuki, and Hubbard 2021; Monti et al. 2022). In particular, video-sharing websites like YouTube have emerged as powerful platforms for disseminating information and mobilizing communities (Uldam and Askanius 2013).
@@
In the past decade, social media facilitated the rapid emergence of social movements that weave narratives in favor of environmental protection and climate action. Research suggests that climate discussions on social platforms, especially those related to initiatives that focus on daily life activities, can significantly enhance public awareness of environmental issues (Mavrodieva et al. 2019). One such activity is the transition to a plant-based diet, which is seen as a tangible and effective action to mitigate the environmental damage caused by over-consumption of animal-based products (Masson-Delmotte et al. 2019; Judge, Fernando, and Begeny 2022). Over the past decade, pro-vegan and pro-vegetarian movements have launched various challenges to raise awareness and encourage the public’s transition towards a plant-based diet. A prominent example is Veganuary, an initiative started in 2014 which encourages people to embrace the vegan diet for the entire month of January each year. In 2023, the initiative attracted over 700k official participants worldwide (veganuary.com 2023). The two other largest initiatives of this kind are Meatless March and No Meat May. The success of these challenges relies heavily on participants and activists sharing their experiences online, as this generally fosters a sense of community and increases engagement (Hou 2023).
@@
The effectiveness of these pro-environmental campaigns is thought to be closely linked to the types of narratives that they employ. The framing of environmental issues can influence attention and shape participation in political debates (Nisbet and Newman 2015), for instance impacting the reactions to environmental scandals (Torelli, Balluchi, and Lazzini 2020). Incorporating persuasive communication strategies into pro-environmental messages can shape the public perceptions on climate change and facilitate behavioral shifts (Pelletier and Sharp 2008). However, despite the extensive literature on storytelling in social movements (Davis 2002; Fiskio 2012; Napoli and Ouschan 2020) and the widely reported presence of collective action in climate change discourse (Bamberg, Rees, and Seebauer 2015; Van Zomeren, Pauls, and Cohen-Chen 2019; Schmitt et al. 2019; Hamann and Reese 2020; Furlong and Vignoles 2021; Gulliver, Fielding, and Louis 2021; Judge, Fernando, and Begeny 2022; Suitner et al. 2023), there is very little quantitative evidence of how much different narratives on climate change discussed in theoretical research are effectively used in social media, and how strongly they are associated with a public response that pledges support to climate action. Addressing this gap is crucial to assess the applicability of existing theoretical frameworks to common social media practices and to enhance our understanding of the persuasive potential of different narrative types.
@@
In an effort to bridge this knowledge gap, we deleved into an exploration of vegan movements and their social campaigns as a form of pro-environmental action. We operationalized a state-of-the-art theoretical framework of narratives that classifies stories pertinent to the vegan ideology using moral foundations (Napoli and Ouschan 2020). We applied this framework to YouTube videos featuring plant-based challenges (such as Veganuary, Meatless March, and No Meat May) to categorize the narratives they present. We did so to answer the following research questions:
@@
Figure 1: Methodological framework.
Our findings indicate that videos whose narratives promote social fight, either in form of protesting or attempting to convert others to the cause, tended to attract reactions characterized by a greater level of collective action. These videos, along with those emphasizing freedom of choice, demonstrated the highest levels of semantic consistency across creators, and the highest semantic alignment between the video transcript and the respective comments. Overall, we found that increased frequency of mentions of collective action in video reactions was significantly associated with a smaller size of the crowd of commenters, a greater content coherence within the reference narrative cluster, and higher frequency of linguistic markers expressing loyalty, a moral dimension that characterizes narratives about social fight, freedom of choice, and duty to educate others through inspirational examples.
@@
Given our particular interest in plant-based diets as a response to climate change, we narrow our focus to narratives within the vegan movement. Specifically, we adopt the narrative framework proposed by Napoli and Ouschan (2020), which focuses on the vegan movement as a specific case of activism, and describes archetypes of storytelling within the vegan ideology. While such a framework predominantly emphasizes animal welfare over explicit climate change concerns, its broad scope allows for a versatile shift towards generic environmental protection without sacrificing relevance. Due to its semantic completeness, this framework offers a more precise description of narratives compared to both general social movement frameworks and other vegan storytelling definitions. In the remainder of the paper, we will refer to the framework as the moral vegan ideology. Next, we will describe the framework itself, outline its theoretical underpinnings, and propose a way to operationalize it.
@@
Table 1: Six narratives types introduced by Napoli and Ouschan (2020), defined as a combination of identity expression (columns) and moral foundations (rows). Each narrative is divided into sub-categories shown in parenthesis. Sub-categories highlighted in bold are those that we found empirically in YouTube videos.
Veganism narratives
@@
where $\emph{f}_{\bullet}$ denotes the relative frequency of a pronoun group in the text. Values around 0.5 suggest a balance between the use of first-person singular and plural pronouns, while values closer to 0 or 1 indicate a focus on the community or the self, respectively. Values shifted toward the two extremes of the range denote “communal-oriented” ($index\leq 0.4$) vs. “agency-oriented” ($index\geq 0.6$) narratives.
@@
We used the YouTube API to analyze discussions related to three plant-based challenges gathering about 800k officially registered participants worldwide: Veganuary, Meatless March, and No Meat May. Our data retrieval spanned videos posted on the platform during the three months centered on the reference month for each challenge, covering a 10-year period from December 2013 to June 2023. We started the collection process by considering the lower-cased names of the movements as a primary keyword (i.e., veganuary, meatless march, no meat may) and performed an expansion taking the relevant hashtags that were most often co-occurring with them. The final sets of keywords are described in Table 2.
With the exclusion of music videos (YouTube video category number 10), which we discarded, we retrieved a total of 12,753 contributions with more than 5M comments. Appendix A provides a yearly breakdown of the data volume for the data of each challenge.
To analyze video content, we leveraged video transcripts. A subset of videos had the auto-caption option enabled by their creators and thus we were able to retrieve their transcript through the Youtube Transcript package (Depoix 2023). To complete the dataset, we employed the Whisper audio-to-text converter to obtain the missing data (Radford et al. 2023). This tool has demonstrated performance comparable to humans in terms of Word Error Rate in long-form transcription tasks. Our pre-processing steps were kept minimal, involving the exclusion of transcripts with fewer than five unique words and the removal of text marked with music tags indicating musical pieces within a YouTube transcript. Additionally, we filtered the pre-processed set of videos to retain only English content, achieved through the langdetect tool for language detection transposed from Java in Python (Shuyo 2010). Those filters left us with a total of 10,617 videos.
To ensure the selection of pertinent content that could effectively be mapped to theory-defined narratives and evoke specific collective reactions, we employed Latent Dirichlet Allocation (LDA) (Blei, Ng, and Jordan 2003) for topic modeling and conducted a manual inspection of the results. To weed out irrelevant content from the rest, we ran LDA with two as the number of topics, and observed that one predominantly revolved around recipes and ingredients, while the other delved into plant-based challenges and their associated ethical aspects. An exam of a sample of the most representative documents for each topic revealed that content within the first group, constituting the majority of data, focused merely on describing recipes and thus showed no relevance to the moral vegan ideology. Since we privilege high precision rather than high recall in our mapping approach, we filtered out those videos reaching a total volume of 3,547 data points, 1,801 of which have at least one comment with more than five unique words.
For each of the three challenges, we gathered a baseline set of videos. These videos served us as a benchmark for re-scaling the metrics calculated on the target video set. To select the baseline, we identified the most frequent video category for each challenge. We then sampled a set of videos of a size similar to the target set (see Appendix A) and within the same reference timeframe, without applying any keyword filter.
Last, we gathered the comments corresponding to our final set of videos. Specifically, we used the CommentThreads method in the YouTube API to gather up to 1000 comments and first-level replies for each video. In total, we collected 516,207 comments on YouTube, which were further refined to 416,736 after simple pre-processing (removing mentions and URLs, and filtering out comments with fewer than five unique words).
@@
Table 2: Keywords for data collection by challenge.
@@
To map videos to narrative types defined by the moral vegan ideology, we first grouped content by measuring the level of collective identity expressed in the transcripts (Eq. 1). We grouped videos into communal-oriented (CI index $\leq 0.4$) and agency-oriented (CI index $\geq 0.6$). We discarded the set of videos with CI index scores between 0.4 and 0.6, as they could not be strongly characterized according to this theoretical dimension. We applied mformer to the remaining 3,045 videos to extract their moral foundation scores. To find clusters of narratives types, we used UMAP (McInnes, Healy, and Melville 2018) to reduce the scores to a 2-dimensional space, and finally applied the HDBSCAN clustering algorithm (Campello, Moulavi, and Sander 2013).
@@
We validated the resulting video content clusters through a double annotation step, considering LLM-based annotation and manual inspection. Given that open-source models, such as Llama- 2 (Touvron et al. 2023), have proven useful assistants to human annotators within the context of multi-class classification of social science concepts (Ziems et al. 2023), we relied on Llama-2 70B Chat for the task. For each cluster in the communal-oriented and agency-oriented groups, we selected the top-30 videos closest to the cluster centroid and we prompted the Llama-2 model to label each transcript according to the most relevant narrative type, if any. We did so by providing a definition of narrative types for each identity group in the LLM prompt, giving the instruction to output other in the case of uncertainty (see Appendix B for the description of prompts). For each video within an HDBSCAN-defined cluster, we thus obtained a label from the LLM. To assign an overall label to the cluster, we considered the percentage of each label within each narrative cluster. We observed a tendency of the LLM to predominantly output one single label per collective identity group (i.e., “duty to educate” in the communal-oriented group and “right to good health” in the agency-oriented one), likely influenced by its internal knowledge bias. To correct for such a bias and explore the nuanced semantic differences between two or more clusters receiving the same overall label, we manually inspected video transcripts to define the theme characterizing each narrative group.
@@
To further enrich our understanding of collective action markers, we incorporated two additional semantic similarity measures that previous research has associated with collective action. First, we measured the semantic coherence of videos within the same narrative group, as consistent narratives can better foster a shared identity among community members, facilitating collective action efforts (Blair and Jost 2003; Van Zomeren, Postmes, and Spears 2008). Second, we measured the semantic alignment between the content of a video and its corresponding comments as a proxy for agreement between influential leaders and activists, which is typically associated with successful collective movements (Thomas, McGarty, and Mavor 2009). Last, we fit an Ordinary Least Squares (OLS) regression to determine whether the variability in collective action responses to videos is significantly associated with the moral features emerging from the video, semantic coherence with videos’ reference narrative, and video-comments alignment.
@@
Content coherence
To gauge the semantic consistency of a narrative type, we evaluated the silhouette score for each video within it, using S-BERT embeddings (Reimers and Gurevych 2019) (all-MiniLM-L6-v2 model) of video transcripts. The silhouette score is a widely-used metric to evaluate cluster quality, and measures the similarity of a datapoint with others in its own cluster compared to the similarity with points in other clusters. The distribution of this silhouette scores across videos associated with a given narrative type assesses the content coherence of such videos in terms of their placement in the specific narrative cluster.
Narrative-reaction alignment
[…]
@@ Results
Using the Collective Identity index (Eq. 1), we identified 389 communal and 2,656 agency-oriented videos. Within each group, the clustering on the UMAP-reduced dimensions of the Moral Foundations Theory (MFT) resulted in four clusters, as shown in Figure 2. We aligned each cluster with theory-defined narrative types through both manual assessment and Llama annotations. Table 3 summarizes the labeling of narratives and provides a video transcript example for each of them. We also characterized each narrative based on the distribution of MFT scores of their respective videos (Figure 3).
@@
We examined the frequency of words associated with collective action in comments on videos across narrative groups (Figure 4). We observed a significant variation in the frequency of these linguistic markers across narratives, with the highest frequency in videos that fell under the fight (convert) narrative. Appendix C reports some examples of comments characterized by high levels of collective action.
@@
The observed macro-level association between collective action indicators, narrative types, semantic coherence of content, and alignment of comments to videos, raises the question of whether the frequency of collective action mentions in comments on individual videos can be inferred from these factors. To address this, we employed an Ordinary Least Squares (OLS) model, using individual videos as the unit of analysis, to predict the degree of commitment based on moral foundations scores, semantic coherence between videos and narrative groups, and video-comment alignment. We also factored in the number of comments on each video as a control variable to account for the video’s popularity and the content creator’s influence. The model was applied after adjusting skewed predictors and calculating the z-scores of all variables, as summarized in Table 4. Despite the inherent challenge in predicting comment commitment variability at the individual video level, we identified several regression coefficients of statistical significance. Specifically, a smaller audience size, higher semantic coherence between content and narrative type, and a discourse more focused on group self-sacrifice (key characteristics of the loyalty moral dimension) were significantly correlated with increased collective action in responses.
@@
While our study provides novel quantitative insights into the relationship between narratives and collective action, it has limitations that open up avenues for future research. Notably, our data pre-processing pipeline retained only 23% of the initial 10,617 videos potentially relevant to the veganism discourse. This was due to several filtering steps aimed at prioritizing precision over recall. Most notably, we excluded video-recipes and all videos not heavily characterized on the Collective Identity axis of our reference theory. Our operationalization covered only a subset of the narratives originally formulated by the theory, which could be due to these filtering steps or to the limited applicability of the theory to the specific context of YouTube. Another limitation pertains to the complex nature of the concepts characterizing the narratives and the public response they received. Despite using validated tools, the inherent ambiguity and complexity of some of the key theoretical notions we base our work on may have introduced noise in the final mapping between empirical clusters of videos and theoretical narratives. Last, the observed collective action in response to narrative content does not necessarily translate to offline participation in social movements, as social media often fosters a form of activism known as “slacktivism” (Morozov 2009).
@@
To protect individual privacy, our dataset includes only aggregated measures for comments, thereby preventing the identification of individual users.

### v2

@@ Introduction
Storytelling is a cornerstone of modern communication, with its influence permeating diverse realms from marketing to political discourse (Lund, Cohen, and Scarles 2018; Seargeant 2020). The power of compelling narratives extends beyond individual persuasion and holds the potential to catalyze collective shifts in opinion, mobilize consensus, and foster social cooperation toward shared goals. Social media have democratized opportunities for creating and sharing narratives on a global scale, empowering individuals to shape opinion formation, democratic deliberation, and collective action (Yasseri et al. 2016; Jennings, Suzuki, and Hubbard 2021; Monti et al. 2022). In particular, video-sharing websites like YouTube have emerged as powerful platforms for disseminating information and mobilizing communities (Uldam and Askanius 2013).
@@
In the past decade, social media facilitated the rapid emergence of social movements that weave narratives in favor of environmental protection and climate action. Research suggests that climate discussions on social platforms, especially those related to initiatives that focus on daily life activities, can significantly enhance public awareness of environmental issues (Mavrodieva et al. 2019). One such activity is the transition to a plant-based diet, which is seen as a tangible and effective action to mitigate the environmental damage caused by over-consumption of animal-based products (Masson-Delmotte et al. 2019; Judge, Fernando, and Begeny 2022). Over the past decade, pro-vegan and pro-vegetarian movements have launched various challenges to raise awareness and encourage the public’s transition towards a plant-based diet. A prominent example is Veganuary, an initiative started in 2014 encouraging people to embrace the vegan diet for the entire month of January each year. In 2023, the initiative attracted over 700k official participants worldwide (veganuary.com 2023). The two other largest initiatives of this kind, spanning over a month each year, are Meatless March and No Meat May. The success of these challenges relies heavily on participants and activists sharing their experiences online, as this generally fosters a sense of community and increases engagement (Hou 2023).
@@
The effectiveness of these pro-environmental campaigns is thought to be closely linked to the types of narratives that they employ (Fernandez et al. 2015). The framing of environmental issues can influence attention and shape participation in political debates (Nisbet and Newman 2015), for instance impacting the reactions to environmental scandals (Torelli, Balluchi, and Lazzini 2020). Incorporating persuasive communication strategies into pro-environmental messages can shape the public perceptions of climate change and facilitate behavioral shifts (Pelletier and Sharp 2008). However, despite the extensive literature on storytelling in social movements (Davis 2002; Fiskio 2012; Napoli and Ouschan 2020) and the widely reported presence of collective action in climate change discourse (Bamberg, Rees, and Seebauer 2015; Van Zomeren, Pauls, and Cohen-Chen 2019; Schmitt et al. 2019; Hamann and Reese 2020; Furlong and Vignoles 2021; Gulliver, Fielding, and Louis 2021; Judge, Fernando, and Begeny 2022; Suitner et al. 2023), there is very little quantitative evidence of how much different narratives on climate change covered by theoretical research are effectively used in social media, and how strongly they are associated with a public response that pledges support to climate action. Addressing this gap is crucial to assess the applicability of existing theoretical frameworks to common social media practices and to enhance our understanding of the persuasive potential of different narrative types.
@@
In an effort to bridge this knowledge gap, we delved into an exploration of vegan movements and their social campaigns as a form of pro-environmental action. We operationalized a state-of-the-art theoretical framework of narratives that classifies stories pertinent to the vegan ideology using moral foundations (Napoli and Ouschan 2020). We applied this framework to YouTube videos featuring plant-based transition challenges (such as Veganuary, Meatless March, and No Meat May) to categorize the narratives they present. We did so to answer the following research questions:
@@
Our findings indicate that videos whose narratives promote social fight, either in the form of protesting or attempting to convert others to the cause, tended to attract reactions characterized by a more frequent presence of linguistic markers that hint at collective action. These videos, along with those emphasizing freedom of choice, demonstrated the highest levels of semantic consistency across creators, and the highest semantic alignment between the video transcript and the respective comments. Overall, we found that an increased frequency of mentions of collective action in video reactions was significantly associated with a smaller size of the crowd of commenters, a greater content coherence of a video within the reference narrative cluster, and a higher frequency of linguistic markers expressing loyalty, a moral dimension that characterizes narratives about social fight, freedom of choice, and duty to educate others through inspirational examples.
@@
Given our specific interest in plant-based diets as a response to climate change, we narrow our focus to narratives within the vegan movement. Within this narrowed scope, scholars have primarily relied on qualitative studies and interviews to infer common narrative patterns. Through an in-depth analysis of vegan blogs, Napoli and Ouschan (2020) found six narrative types that are described with a combination of moral foundations and identity framing that emerge from the stories. Compared to other theoretical categorizations of vegan narratives types (Waters 2022; Aavik 2023), the framework by Napoli and Ouschan (2020) is semantically more comprehensive, and offers an operationalizable map within the context of online content. Therefore, we adopt such a narrative framework as our theoretical reference, focusing on the vegan movement as a specific case of activism. While the framework predominantly emphasizes animal welfare over explicit climate change concerns, its broad scope allows for a versatile shift towards generic environmental protection without sacrificing relevance. In the remainder of the paper, we will refer to the framework as the moral vegan ideology. Next, we will describe the framework itself, outline its theoretical underpinnings, and propose a way to operationalize it.
Veganism narratives
@@
Table 1: Six narratives types introduced by Napoli and Ouschan (2020), defined as a combination of identity expression (columns) and moral foundations (rows). Each narrative is divided into sub-categories shown in parenthesis. Sub-categories highlighted in bold are those that we found empirically in YouTube videos through our experiments.
@@
where $\emph{f}_{\bullet}$ denotes the relative frequency of a pronoun group in the text. The index is centered around 0.5: values around 0.5 suggest a balance between the use of first-person singular and plural pronouns, while values closer to 0 or 1 indicate a focus on the community or the self, respectively. Values shifted toward the two extremes of the range denote “communal-oriented” ($index\leq 0.4$) vs. “agency-oriented” ($index\geq 0.6$) narratives.
@@
We used the YouTube API to analyze discussions related to three plant-based challenges gathering about 800k officially registered participants worldwide: Veganuary, Meatless March, and No Meat May. Our selection criteria were driven by the widespread participation and popularity of Veganuary, the leading challenge in this domain. In addition, we considered Meatless March, and No Meat May as they are analogous time-constrained initiatives that enjoy a substantial level of public interest. Our data retrieval spanned videos posted on the platform during the three months centered on the reference month for each challenge, covering a 10-year period from December 2013 to June 2023. We started the collection process by considering the lower-cased names of the movements as a primary keyword (i.e., veganuary, meatless march, no meat may) and performed an expansion taking the relevant hashtags that were most often co-occurring with them. In practice, we first queried the API using the seed keywords, extracted the ten most frequent keywords in the descriptions of the set of retrieved videos, and manually selected those that were semantically relevant to the challenge. The final sets of keywords are in Table 2.
@@
Table 2: Keywords used for the data collection by initiative.
We retrieved a total of 12,753 contributions with more than 5M comments. Appendix A provides a yearly breakdown of the data volume for the data of each challenge.
To analyze video content, we leveraged video transcripts. We filtered the set of videos to retain only English content, achieved through the langdetect tool for language detection transposed from Java in Python (Shuyo 2010). A subset of videos had the auto-caption option enabled by their creators and thus we were able to retrieve their transcript through the Youtube Transcript package (Depoix 2023). To complete the dataset, we employed the Whisper audio-to-text converter to obtain the missing data (Radford et al. 2023). This tool has demonstrated performance comparable to humans in terms of Word Error Rate in long-form transcription tasks.
To ensure the selection of pertinent content that could effectively be mapped to theory-defined narratives and evoke specific collective reactions, we employed Latent Dirichlet Allocation (LDA) (Blei, Ng, and Jordan 2003) for topic modeling and conducted a manual inspection of the results. To weed out irrelevant content from the rest, we ran LDA with two as the number of topics, we identified the ten most common words per topic and examined the top-5 most representative video transcripts for each topic. We observed that one topic predominantly revolved around recipes and ingredients, showing no relevance to the moral vegan ideology theory, while the other delved into plant-based challenges and their associated ethical aspects. In our narrative mapping approach, the emphasis lies on achieving high precision rather than high recall. This entails prioritizing the accurate identification of pertinent elements, even at the expense of potentially overlooking some less relevant ones. As such, we filtered out recipe-focused videos (i.e. one of the retrieved topics) reaching a total volume of 3,547 data points, 1,801 of which have at least one comment with more than five unique words.
Generally speaking, our pre-processing steps were kept minimal, involving the exclusion of transcripts with fewer than five unique words and the removal of text marked with music tags indicating musical pieces within a YouTube transcript. Table 3 summarizes the filtering steps and resulting data volumes.
Original English
With
transcript
Valid
topic
Valid
comments
12,753 11,854
8,954
3,547 1,801
2,399
Table 3: Video data volume by filtering step. Filters are applied in series, left to right.
For each of the three challenges, we gathered a baseline set of videos. These videos served us as a benchmark for re-scaling the metrics calculated on the target video set. To select the baseline, we identified the most frequent YouTube video category for each challenge. Within such categories, we then collected a set of videos of a size similar to the target set (see Appendix A) and within the same reference timeframe, without applying any keyword filter.
Last, we gathered the comments corresponding to our final set of videos. Specifically, we used the CommentThreads method in the YouTube API to gather up to 1000 comments and first-level replies for each video. We collected the most recent comments because, at the time of writing, the YouTube API does not allow for retrieving a random sample of comments. In total, we collected 516,207 comments on YouTube, which were further refined to 416,736 after simple pre-processing (removing mentions and URLs, and filtering out comments with fewer than five unique words).
@@
To map videos to narrative types defined by the moral vegan ideology, we first grouped content by measuring the level of collective identity expressed in the transcripts (Eq. 1). We grouped videos into communal-oriented (CI index $\leq 0.4$) and agency-oriented (CI index $\geq 0.6$). We discarded the set of videos with CI index scores between 0.4 and 0.6, as they could not be strongly characterized according to this theoretical dimension. We applied mformer to the remaining 3,045 videos to extract their moral foundation scores. To find clusters of narrative types, we used UMAP (McInnes, Healy, and Melville 2018) to reduce the scores to a 2-dimensional space, and finally applied the HDBSCAN clustering algorithm (Campello, Moulavi, and Sander 2013).
@@
We employed manual annotation to validate the resulting video content clusters. Acknowledging the utility of open-source models like Llama-2 (Touvron et al. 2023) in aiding human annotators in the classification of social science concepts (Ziems et al. 2023), we considered Llama-2 70B Chat as an initial supporting tool. For each cluster in the communal-oriented and agency-oriented groups, we selected the top 30 videos closest to the cluster centroid and prompted the language model to label each transcript within this subset according to the most relevant narrative type, if discernible. We defined narrative types for each identity group in the prompt and instructed it to output “other” in cases of uncertainty (refer to Appendix B for prompt descriptions). Notably, we observed a tendency of the LLM to predominantly output a single label per collective identity group (e.g., “duty to educate” in the communal-oriented group and “right to good health” in the agency-oriented one), likely influenced by its internal knowledge bias. Given such a first initial guideline, we engaged in manual inspection of transcripts within the top-30 videos subset. This step allowed us to label each cluster and qualitatively define the theme characterizing each narrative group based on sub-categories extracted from the theory, described in Table 1. Data volume for each annotated narrative group can be found in Appendix A.
@@
To further enrich our understanding of collective action markers, we incorporated two additional semantic similarity measures that previous research has associated with collective action. First, we measured the semantic coherence of videos within a narrative group, as consistent narratives can better foster a shared identity among community members, facilitating collective action efforts (Blair and Jost 2003; Van Zomeren, Postmes, and Spears 2008). Second, we measured the semantic alignment between the content of a video and its corresponding comments as a proxy for agreement between influential leaders and activists, which is typically associated with successful collective movements (Thomas, McGarty, and Mavor 2009). Last, we fit an Ordinary Least Squares (OLS) regression to determine whether the variability in collective action responses to videos is significantly associated with the moral features emerging from them, the semantic coherence with videos’ reference narratives, and the video-comments alignment.
@@
Content coherence
To gauge the semantic consistency of a narrative type, we evaluated the silhouette score for each video within it, using S-BERT embeddings (Reimers and Gurevych 2019) (all-MiniLM-L6-v2 model) of video transcripts. The embedding model, trained on sentence-level data, demonstrates strong performance across various natural language processing downstream tasks. Nonetheless, its effectiveness may fluctuate when handling very large documents, often due to truncation issues. The silhouette score is a widely-used metric to evaluate cluster quality, and measures the similarity of a datapoint with others in its own cluster compared to the similarity with points in other clusters. The distribution of silhouette scores across videos associated with a given narrative type assesses the content coherence of such videos in terms of their placement in the specific narrative cluster.
Narrative-reaction alignment
[…]
Results
@@
Using the Collective Identity index (Eq. 1), we identified 389 communal and 2,656 agency-oriented videos. Within each group, the clustering on the UMAP-reduced dimensions of the Moral Foundations Theory (MFT) resulted in four clusters, as shown in Figure 2. We aligned each cluster with theory-defined narrative types through manual assessment. Table 4 summarizes the labeling of narratives and provides a video transcript example for each of them. We also characterized each narrative based on the distribution of MFT scores of their respective videos (Figure 3).
@@
We examined the frequency of words associated with collective action in comments on videos across narrative groups (Figure 4). We observed a variation in the frequency of these linguistic markers across narratives, with the highest frequency in videos that fell under the fight (convert) narrative. Appendix C reports some examples of comments characterized by high levels of collective action.
@@
The observed macro-level association between collective action indicators, narrative types, semantic coherence of content, and alignment of comments to videos, raises the question of whether the frequency of collective action mentions in comments on individual videos can be inferred from these factors. To address this, we employed an Ordinary Least Squares (OLS) model, using individual videos as the unit of analysis, to predict the degree of commitment based on moral foundations scores, semantic coherence between videos and narrative groups, and video-comment alignment. We also factored in the number of comments on each video as a control variable to account for the video’s popularity and the content creator’s influence. The model was applied after adjusting skewed predictors and calculating the z-scores of all variables, as summarized in Table 5. Despite the inherent challenge in predicting comment commitment variability at the individual video level, we identified several regression coefficients of statistical significance (p-values $\leq 0.1$). Specifically, a smaller audience size, higher semantic coherence between content and narrative type, and a discourse more focused on group self-sacrifice (key characteristics of the loyalty moral dimension) were significantly correlated with increased collective action in responses.
@@
While our study provides novel quantitative insights into the relationship between narratives and collective action, it has limitations that open up avenues for future research. Primarily, the vegan narrative framework we employed is inherently tailored to its specific topic, as defined within the existing literature. Nonetheless, its structured and clear outline, rooted in collective identity and moral perspectives, presents the opportunity for expansion into other contexts involving lifestyle choices. Other issues arise from the limited amount of data involved in the final steps of the study. Notably, our data pre-processing pipeline retained only 24% of the initial 12,753 videos potentially relevant to the veganism discourse. This was due to several filtering steps aimed at prioritizing precision over recall. Most notably, we excluded video-recipes and all videos not heavily characterized on the Collective Identity axis of our reference theory. Our operationalization covered only a subset of the narratives originally formulated by the theory, which could be due to these filtering steps or to the limited applicability of the theory to the specific context of YouTube. Another limitation pertains to the complex nature of the concepts characterizing the narratives and the public response they received. Despite using validated tools, the inherent ambiguity and complexity of some of the key theoretical notions we base our work on may have introduced noise in the final mapping between empirical clusters of videos and theoretical narratives. Last, the observed collective action in response to narrative content does not necessarily translate to offline participation in social movements, as social media often fosters a form of activism known as “slacktivism” (Morozov 2009).
Despite these constraints, our study successfully captures the existence of nuances in pro-environmental storytelling and discerns variations in the type of engagement that different narrative types receive.
@@
The dataset associated with this paper complies with the FAIR principles (FORCE11 2020), as outlined in the DataSheet (Supplementary Material).
Code and data availability
All code used for the experiments and information about the videos examined, in terms of narrative orientations, statistics, metrics specific to this paper, reference baseline videos, and comments retrieved and analyzed is available on a GitHub repository [^1] https://github.com/ariannap13/VeganCollectiveAction.
Acknowledgments
We acknowledge the support from the Carlsberg Foundation through the COCOONS project (CF21-0432). The funder had no role in study design, data collection and analysis, decision to publish, or preparation of the manuscript.